OpenStudy (anonymous):

The base of an isosceles triangle is half as long as the two equal sides. Write the area of the triangle as a function of the length of the base

OpenStudy (anonymous):

A = 1/2 b*h, is the original formula. In your case, you need to separate the variables. A = [(1/2*b)*h] For example, say 3 is the base and 4 is the height. Original Formula: A = 1/2 3*4 = 6 Your Problem Formula: A= [(1/2*3) *4] = 1.5*4 = 6 So your answer is A = [(1/2*b)*h]

OpenStudy (anonymous):

I understand it more now but why would I have to separate the variables? Thanks so much for the help!

OpenStudy (anonymous):

You have to separate the variables because that is what the question is asking. The base is half as long as the two equal sides, so you need to illustrate that in equation form. It's just a conversion. If it said the base was a fourth as long, you would do A = 1/2 *1/4 of the base. It's just what the question is asking you to modify into your new formula.

OpenStudy (anonymous):

Oh okay I understand. Thanks so much!